Find the points of intersection of the circle with the line with the equation given.
(x+ 6 )^2 + ( y - 3)^2 = 24: y = x + 2

Answer by ikleyn(52824)

.
Find the points of intersection of the circle with the line with the equation given.
(x+ 6 )^2 + ( y - 3)^2 = 24: y = x + 2
-------------------------------------------------

Substitute this expression for y: y = x + 2 into the equation of the circle.

You will get the quadratic equation for x.

Solve it and find the roots (x's).

Then restore the corresponding y's.